OpenStudy (animelover91804):

The total snowfall in Burke, Virginia, this January was 7 1/4 in. In February, 20% more snow fell than in January. How much snow fell in February in Burke, Virginia? Express your answer as a mixed number in simplest form.

OpenStudy (anonymous):

it is 85,2/4

OpenStudy (anonymous):

we firnd how much snowfalls is 20% 71/4*0.2=14.2/4 next we find how much was in february 71/4+14,2/4=85,2/4